Is energy produced when a transformer steps up the voltage ?

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Nikhil garg

NO,energy does not produce, it remains constant

for transformer power=V*I=constant

                 V1*I1=V2*I2=constant

for steps up voltage only voltage increase & cureent decrease, while power remains same.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.


AKASH GOYAL
AskiitiansExpert-IITD